What is the average density of a woman who floats in freshwater with 3% of her volume above the surface?


1
Expert's answer
2022-04-27T13:34:25-0400
"mg=g\rho_wV_s"

"g\rho V=g\rho_wV_s"

"\rho=\rho_w\frac{V_s}{V}=1000*0.97=970\:\rm kg/m^3"